The pH of a 0.001M HCl solution is ___. The pH of a sample of milk is 6.5, while a sample of a soft drink has a pH of 4.5. The hydrogen ion concentrations (H+) of the two samples are:

A. 2; 2 times different

B. 2; 100 times different

C. 3; 2 times different

D. 3; 100 times different
Answer #1

Answer

D)

3; 100times different

Explanation

i) HCl is strong acid , so

[HCl] = [H+]

pH = -log[H+]

pH = -log(0.001M)

pH = 3

ii)

Milk

-log[H+] = 6.5

[H+] = 1×10-6.5M

[H+] = 3.16 ×10-7M

Soft drink

-log[H+] = 4.5

[H+] = 1×10-4.5

[H+] = 3.16 ×10-5

Therefore,

H+ concentration in soft drink is 100time higher than H+ concentration in milk.